Phone calls enter the ”support desk” of an electricity supplying company on the

average two every 3 minutes. If one assumes an approximate Poisson process:

(i) What is the probability of no calls in 3 minutes?

[2 marks]

(ii) What is the probability of utmost 6 calls in a 9 minute period?


1
Expert's answer
2021-09-08T07:08:05-0400

Let "X=" the number of calls: "X\sim Po(\lambda t)."

(i)


"\lambda t=2"

"P(X=0)=\dfrac{e^{-\lambda t}(\lambda t)^0}{0!}=e^{-2}\approx0.135335"

(ii)


"\lambda t=6"

"P(X\leq6)=P(X=0)+P(X=1)+P(X=2)"

"+P(X=3)+P(X=4)+P(X=5)+P(X=6)"

"=\dfrac{e^{-\lambda t}(\lambda t)^0}{0!}+\dfrac{e^{-\lambda t}(\lambda t)^1}{1!}+\dfrac{e^{-\lambda t}(\lambda t)^2}{2!}+\dfrac{e^{-\lambda t}(\lambda t)^3}{3!}"

"+\dfrac{e^{-\lambda t}(\lambda t)^4}{4!}+\dfrac{e^{-\lambda t}(\lambda t)^5}{5!}+\dfrac{e^{-\lambda t}(\lambda t)^6}{6!}"

"=e^{-6}(1+6+18+36+54+64.8+64.8)"

"=224.6e^{-6}\approx0.606303"